What is the mass of an object with a specific heat capacity of
750J/(kg


C)
750J/(kg∘C)
that needs
5625J
5625J
to have its temperature raised by
15.0


C
15.0∘C
?
1
Expert's answer
2018-06-06T10:43:08-0400
We can find the mass of the object from the formula:
Q=mc∆t,
here, Q=5625 J is the amount of heat that required to raise the temperature of the object by ∆t=15.0 ℃, m is the mass of the object, c=750 J⁄(kg℃) is the specific heat capacity of the object.
Then, we get:
m=Q/c∆t=(5625 J)/(750 J/(kg℃)∙15.0 ℃)=0.5 kg.